complete the square x^2+12x=9
Answer by nerdybill(7384)

complete the square
x^2+12x=9
take half 'b', square it and add to both sides:
x^2+12x+36 = 9+36
x^2+12x+36 = 45
(x+6)(x+6) = 45
(x+6)^2 = 45
x+6 = sqrt(45)
x+6 = +-sqrt(3*3*5)
x+6 = +-3sqrt(5)
x = -6+-3sqrt(5)